# From Raw Corpora to Domain Benchmarks: Automated Evaluation of LLM Domain Expertise

### v2

@@ Abstract
Accurate domain-specific benchmarking of LLMs is essential, specifically in domains with direct implications for humans, such as law, healthcare, and education. However, existing benchmarks are documented to be contaminated and are based on multiple choice questions, which suffer from inherent biases. To measure domain-specific knowledge in LLMs, we present a deterministic pipeline that transforms raw domain corpora into completion-style benchmarks without relying on other LLMs or costly human annotation. Our approach first extracts domain-specific keywords and related target vocabulary from an input corpus. It then constructs prompt-target pairs where domain-specific words serve as prediction targets. By measuring LLMs’ ability to complete these prompts, we provide a direct assessment of domain knowledge at low computational cost. Our pipeline avoids benchmark contamination, enables automated updates with new domain data, and facilitates fair comparisons between base and instruction-tuned (chat) models.
We validate our approach by showing that model performances on our benchmark significantly correlate with those on an expert-curated benchmark. We then demonstrate how our benchmark provides insights into knowledge acquisition in domain-adaptive, continual, and general pretraining. Finally, we examine the effects of instruction fine-tuning by comparing base and chat models within our unified evaluation framework. In conclusion, our pipeline enables scalable, domain-specific, LLM-independent, and unbiased evaluation of both base and chat models.
@@ 1 Introduction
Existing approaches measure proxies of domain knowledge rather than domain knowledge itself. Perplexity, the most common metric for comparing language models, aggregates prediction quality over all tokens, domain-relevant and domain-irrelevant alike, making it impossible to disentangle genuine domain knowledge from general linguistic fluency (Öncel et al., 2024; Xu et al., 2024; Meister and Cotterell, 2021). Multiple-choice question (MCQ) benchmarks such as MMLU (Hendrycks et al., 2020) suffer from well-documented biases: Gupta et al. (2024) show that simply reordering answer choices can substantially alter model accuracy, while Chandak et al. (2025) demonstrate that popular MCQ benchmarks can often be answered without even seeing the question, calling into question whether they measure knowledge at all. MCQ formats further disadvantage base models, which lack instruction-following capabilities and are sensitive to few-shot formatting (Brown et al., 2020; Alzahrani et al., 2024). Beyond format-level issues, benchmark contamination, where models are evaluated on data they were trained on, undermines the validity of static benchmarks and is difficult to detect post hoc (Zhou et al., 2023). Finally, existing domain-specific benchmarks such as MedQA (Jin et al., 2021) and LegalQA (Louis et al., 2024) organize around broad disciplinary categories rather than fine-grained subdomains, and rely on formats (MCQs, long-form answers) that make fair comparison across model types difficult.
Figure 1: Issues with existing domain-specific benchmarks: Perplexity aggregates predictions over all tokens (including domain-irrelevant ones); performance on multiple choice questions depend on the order of options; many benchmarks are already incorporated in the training sets of LLMs; and manual creation is simply too expensive.
In this work, we present a deterministic pipeline that transforms any raw domain corpus into a completion-style benchmark with prompt–target pairs, without requiring manual curation or reliance on other LLMs during construction. By measuring a model’s ability to predict domain-specific target terms given contextual prompts, we directly assess domain knowledge through the same next-token prediction objective on which all LLMs, both base and instruction-tuned, are pretrained. We evaluate models by computing the rank of the correct target token in the model’s output distribution, providing a soft and interpretable metric of domain expertise.
@@
Figure 2: A conceptual overview of our proposed pipeline for generating a completion-based benchmark from a raw domain corpus. Our pipeline extracts and refines keywords from the input corpus. It then matches each sentence with relevant keywords to focus evaluation on domain-relevant content. For every keyword, we construct a target vocabulary by collecting domain-specific phrases from the matched sentences; these phrases will serve as prediction targets. The domain expertise is quantified by how well the models complete these targets from given prompts.
@@ 2.1 Data Curation
We select four diverse evaluation domains that span different scientific fields: Computer Science – Artificial Intelligence (CS.AI), Physics and Society (Physics (Soc-Ph)), Quantitative Biology – Populations and Evolution (Q-Bio.PE), and General Economics (Econ-GN). This selection ensures that our validation covers domains with varying corpus sizes, vocabularies, and degrees of overlap with typical pretraining data.
@@ 2.2 Keyword Generation
  • Third, we apply adaptive length-based filtering to retain a proportionally balanced subset of n-grams, adjusting length ratios dynamically across domains to target approximately 300 high-quality keywords per domain. We note that more keywords enable finer domain coverage but require greater computational resources, a trade-off practitioners can adjust based on their specific objectives. Our experiments showed consistent results with 150 and 450 keywords, demonstrating robustness to this parameter choice.
  • Finally, we reduce redundancy by computing pairwise cosine similarities between keywords using the all-MiniLM-L6-v2 (Wang et al., 2020) model from the Sentence Transformer library (Reimers and Gurevych, 2019) and merging semantically similar entries with similarity scores above 0.85. This threshold balances two competing objectives: ensuring sufficiently strong semantic connections to eliminate true redundancy while avoiding overly restrictive matching that would merge distinct but related concepts. Lower thresholds (e.g., 0.7) risk conflating semantically distinct keywords, while higher thresholds (e.g., 0.95) may retain near-duplicate entries that provide minimal additional coverage.
Please see Sec. A.1 for details of each step. Table 4 lists ten representative keywords generated by our pipeline for each of the four domains.
@@ 2.4 TF and TF-IDF Target Vocabulary
Table 1: Example keywords with their TF and TF-IDF target vocabularies (10 terms each) for the CS.AI domain.
@@ 4.2 Quantifying Knowledge Acquisition During General Pretraining
We also evaluated the checkpoints on the CS.AI domain, depicted in Supp. Fig. 16, which exhibits a similar trend, verifying the consistency of our evaluation across downstream domains. We note that the overall ranks are considerably better on CS.AI, indicating the relative ease of the task for the model. We attribute this to relatively abundant training data from this domain, resulting in improved performance.
@@ 4.3 Continually Pretraining Llama2-7B model
Following previous experiments, we select CS.AI as the target domain and continually pretrain Llama2-7B on a sequence of domains from M2D2. To construct a meaningful training order, we randomly sample $5000$ abstracts from the arXiv dataset, embed them using Sentence Transformers, and compute cosine similarity between their embeddings and those of CS.AI domain abstracts. We then sort the M2D2 domains by their similarity to CS.AI. From this list, we uniformly sample 12 domains, sort them by dissimilarity to CS.AI, and use them as the continual pretraining sequence for Llama2-7B (see domain names in Fig. 6).
@@ 4.4 Comparing Base and Chat Models
Our rank analysis also allows us to inspect how alignment impacts model knowledge. Fig. 24 reveals that base models generally outperform their chat counterparts, corroborating findings from Bai et al. (2022) who demonstrated that smaller models experience significant “alignment tax” where performance degrades after instruction tuning. This degradation is particularly pronounced for Llama2-7B and Mistral-7B-v0.3, whose performance drop after instruction tuning across all domains, suggesting significant room for improvement in their alignment pipelines. Moreover, our framework provides a more fine-grained view than the generic evaluations in Bai et al. (2022), revealing that some models like Qwen2-1.5B and Llama3.1-8B exhibit smaller degradations or even occasional improvements, demonstrating that alignment effects are highly heterogeneous across architectures and domains. This suggests that alignment may dilute specialized domain knowledge, as the technical vocabulary required for domain expertise may be weakened during instruction tuning that emphasizes broader linguistic competencies.
@@ 7 Limitations
While our proposed pipeline demonstrates strong empirical results across multiple experimental settings, several limitations warrant discussion. Our validation approach uses benchmarks generated by Claude Sonnet 4 as a reference standard. To validate this choice, we created a manual expert benchmark for the CS.AI domain from "Understanding Deep Learning" Prince (2023), which demonstrated strong correlation with Claude-generated benchmarks (r=0.91, p=0.012), confirming Claude’s reliability for this purpose. However, we acknowledge that extending manual validation to all evaluated domains (Physics, Economics, Biology) was beyond the scope of this work due to resource constraints. The strong correlations we observe between our pipeline and Claude benchmarks on CS.AI domain, coupled with intuitive behavioral patterns in controlled experiments, suggest that this validation generalizes beyond CS.AI.
Our experimental scope encompasses academic texts from arXiv spanning four domains from different science fields, which, while providing clean and well-structured data ideal for methodological validation, represents a focused subset of potential applications. Similarly, computational constraints limited our model coverage to eight base models and six chat variants ranging from 1.5B to 8B parameters, and restricted the evaluations in our domain adaptation and continual learning experiments primarily to Computer Science (Artificial Intelligence). While the evaluation of significantly larger models (e.g., >70B parameters) would provide additional validation, our current model range captures diverse training paradigms and architectural designs. These scope limitations are largely driven by practical considerations, but the consistent patterns we observe across diverse STEM domains and model architectures (GPT2, Llama2, Mistral, Qwen2) at different scales provide evidence for broader applicability.
The evaluation of chat models presents additional challenges since these models are optimized for conversational interactions rather than completion tasks. As we demonstrated in our MMLU analysis, chat models exhibit significant sensitivity to prompt formatting and ordering, which could potentially affect our completion-based evaluation. Nevertheless, completion tasks align more closely with the training objective shared by both base and chat models, and the systematic patterns we observe suggest our method captures meaningful signals about knowledge retention.
These limitations, while acknowledging areas for future expansion, do not diminish the fundamental contribution of providing a scalable, automated approach to domain-specific knowledge evaluation that addresses critical gaps in current LLM assessment methodologies. Moreover, our pipeline’s inherently domain-agnostic design can be extended to other corpus types and domains with only minor preprocessing adjustments, while the core methodology remains unchanged.

### v3

@@ Abstract
Accurate domain-specific benchmarking of LLMs is essential, specifically in domains with direct implications for humans, such as law, healthcare, and education. However, existing benchmarks are documented to be contaminated and are based on multiple-choice questions, which suffer from inherent biases. To measure domain-specific knowledge in LLMs, we present a deterministic pipeline that transforms raw domain corpora into completion-style benchmarks without relying on other LLMs or costly human annotation. Our approach first extracts domain-specific keywords and related target vocabulary from an input corpus. It then constructs prompt-target pairs where domain-specific words serve as prediction targets. By measuring LLMs’ ability to complete these prompts, we provide a direct assessment of domain knowledge at low computational cost. Our pipeline avoids benchmark contamination, enables automated updates with new domain data, and facilitates fair comparisons between base and instruction-tuned (chat) models. We validate our approach by showing that model performances on our benchmark significantly correlate with those on an expert-curated benchmark. We then demonstrate how our benchmark provides insights into knowledge acquisition in domain-adaptive, continual, and general pretraining. Finally, we examine the effects of instruction fine-tuning by comparing base and chat models within our unified evaluation framework. In conclusion, our pipeline enables scalable, domain-specific, LLM-independent, and unbiased evaluation of both base and chat models.
@@ 1 Introduction
Existing approaches measure proxies of domain knowledge rather than domain knowledge itself. Perplexity, the most common metric for comparing language models, aggregates prediction quality over all tokens, domain-relevant and domain-irrelevant alike, making it impossible to disentangle genuine domain knowledge from general linguistic fluency (Öncel et al., 2024; Xu et al., 2024; Meister and Cotterell, 2021). Multiple-choice question (MCQ) benchmarks such as MMLU (Hendrycks et al., 2020) suffer from well-documented biases: Gupta et al. (2024) show that simply reordering answer choices can substantially alter model accuracy, while Chandak et al. (2025) demonstrate that popular MCQ benchmarks can often be answered without even seeing the question, calling into question whether they measure knowledge at all. MCQ formats further disadvantage base models, which lack instruction-following capabilities and are sensitive to few-shot formatting (Brown et al., 2020; Alzahrani et al., 2024). Beyond format-level issues, benchmark contamination, where models are evaluated on data they were trained on, undermines the validity of static benchmarks and is difficult to detect post hoc (Zhou et al., 2023). Finally, existing domain-specific benchmarks such as MedQA (Jin et al., 2021) and LegalQA (Louis et al., 2024) organize around broad categories rather than fine-grained subdomains, and rely on formats (MCQs, long-form answers) that make fair comparison across model types difficult.
Figure 1: Issues with existing domain-specific benchmarks: Perplexity aggregates predictions over all tokens (including domain-irrelevant ones); performance on multiple choice questions depend on the order of options; many benchmarks are already incorporated in the training sets of LLMs; and manual creation is too expensive.
In this work, we present a deterministic pipeline that transforms any raw domain corpus into a completion-style benchmark with prompt–target pairs, without requiring manual curation or reliance on other LLMs. By measuring a model’s ability to predict domain-specific target terms given contextual prompts, we directly assess domain knowledge through the same next-token prediction objective on which all LLMs, both base and instruction-tuned, are pretrained. We evaluate models by computing the rank of the correct target token in the model’s output distribution, providing a soft and interpretable metric of domain expertise.
@@
Figure 2: A conceptual overview of our pipeline for generating a completion-based benchmark from a raw domain corpus. We extract and refine keywords from the input corpus. We then match each sentence with relevant keywords to focus evaluation on domain-relevant content. For every keyword, we construct a target vocabulary by collecting domain-specific phrases from the matched sentences; these phrases serve as prediction targets. The domain expertise is quantified by how well the models predict the targets from given prompts.
@@ 2.1 Data Curation
We select four diverse evaluation domains that span different scientific fields: Computer Science – Artificial Intelligence (CS.AI), Physics and Society (Physics (Soc-Ph)), Quantitative Biology – Populations and Evolution (Q-Bio.PE), and General Economics (Econ-GN). Our validation covers domains with varying corpus sizes, vocabularies, and degrees of overlap with typical pretraining data.
@@ 2.2 Keyword Generation
  • Third, we apply adaptive length-based filtering to retain a proportionally balanced subset of n-grams, adjusting length ratios across domains to target 300 high-quality keywords per domain. We note that more keywords enable finer domain coverage but require greater computational resources, a trade-off practitioners can adjust based on their specific objectives. Our experiments showed consistent results with 150 and 450 keywords, demonstrating robustness to this parameter choice.
  • Finally, we reduce redundancy by computing pairwise cosine similarities between keywords using the all-MiniLM-L6-v2 (Wang et al., 2020) from the Sentence Transformer library (Reimers and Gurevych, 2019) and merging semantically similar entries with similarity scores above 0.85. This threshold balances two competing objectives: ensuring sufficiently strong semantic connections to eliminate true redundancy while avoiding overly restrictive matching that would merge distinct but related concepts. Lower thresholds (e.g., 0.7) risk conflating semantically distinct keywords, while higher thresholds (e.g., 0.95) may retain near-duplicate entries that provide minimal additional coverage.
Please see Sec. A.1 for details of each step. Table 4 lists ten representative keywords generated by our pipeline.
@@ 2.4 TF and TF-IDF Target Vocabulary
Table 1: Example keywords with their TF and TF-IDF target vocabularies for the CS.AI domain.
@@ 4.2 Quantifying Knowledge Acquisition During General Pretraining
We also evaluated the checkpoints on the CS.AI domain. The similar trends depicted in Supp. Fig. 16 verify the consistency of our evaluation across downstream domains. We note that the overall ranks are considerably better on CS.AI, indicating the relative ease of the task for the model. We attribute this to relatively abundant training data from this domain, resulting in improved performance.
@@ 4.3 Continually Pretraining Llama2-7B model
Following previous experiments, we select CS.AI as the target domain and continually pretrain Llama2-7B on a sequence of domains from M2D2. To construct a meaningful training order, we randomly sample $5000$ abstracts from arXiv, embed them using Sentence Transformers, and compute cosine similarity between their embeddings and those of CS.AI domain abstracts. We then sort the M2D2 domains by their similarity to CS.AI. From this list, we uniformly sample 12 domains, sort them by dissimilarity to CS.AI, and use them as the continual pretraining sequence for Llama2-7B (see domain names in Fig. 6).
@@ 4.4 Comparing Base and Chat Models
Our rank analysis also allows us to inspect how alignment impacts model knowledge. Fig. 24 reveals that base models generally outperform their chat counterparts, corroborating findings from Bai et al. (2022) who demonstrated that smaller models experience significant “alignment tax” where performance degrades after instruction tuning. This degradation is particularly pronounced for Llama2-7B and Mistral-7B-v0.3, whose performance drop after instruction tuning across all domains, suggesting significant room for improvement in their alignment pipelines. Moreover, our framework provides a more fine-grained view than the generic evaluations in Bai et al. (2022), revealing that some models like Qwen2-1.5B and Llama3.1-8B exhibit smaller degradations or even occasional improvements, demonstrating that alignment effects are highly heterogeneous across architectures and domains.
@@ 7 Limitations
While our proposed pipeline demonstrates strong empirical results across multiple experimental settings, several limitations require discussion. Our validation approach uses benchmarks generated by Claude Sonnet 4 as a reference standard. Although we validate this choice with a manual expert benchmark for CS.AI (r=0.91, p=0.012), extending this manual validation to all four domains was beyond the scope of this work. The strong pipeline-to-Claude correlations across all domains, combined with intuitive patterns in controlled experiments, suggest that this validation generalizes beyond CS.AI.
Our experimental scope is limited to arXiv academic texts across four STEM domains, and computational constraints restricted model coverage to eight base and six chat variants, ranging from 1.5B to 8B parameters. While evaluation of larger models (e.g., >70B) would provide additional validation, consistent patterns across diverse architectures (GPT2, Llama2, Mistral, Qwen2) support broader applicability.
Finally, chat models are optimized for conversational interaction rather than completion tasks, which could affect our evaluation. However, completion tasks align with the pretraining objective shared by both model types, and the systematic patterns we observe suggest our method captures meaningful signals about knowledge retention across both settings.